22 STUDENTS HAVE CHARTERED A BUS TO TRAVEL TO THE CHAMPIONSHIP GAME AND WILL SPLIT THE COST EQUALLY. aFTER 17 OF THE STUDENTS HAVE TURNED IN THEIR SHARE, $131.75 HAS BEEN COLLECTED. What is the cost of chartering the bus?
Answer by edjones(7311)

Let x be the cost of chartering the bus.
17x/22=131.75
17x=131.75*22
x=(131.75*22)/17
x=$170.50
.
Ed